How do you solve this problem, there are two consecutive odd numbers when multiplied equal the product 323. What are the two numbers?
Answer by ankor@dixie-net.com(12692)

there are two consecutive odd numbers
x, and (x+2)
:
when multiplied equal the product 323.
x*(x+2) = 323
:
What are the two numbers?
x^2 + 2x = 323
x^2 + 2x - 323 = 0; a quadratic equation
Factors to:
(x + 19)(x - 17) = 0
Two solutions:
x = - 19; the two numbers would be -19 and -17
and
x = + 17: the two number would be +17 and + 19
:
Did this make sense to you. Not that hard, right?